Rajesh gave 40% of his accumulated amount to his wife,25% to his son & spent 60% on education in the remaining amount.now he is with 2700.what is the amount he gave to his wife
Answer by Theo(13342)

the amount he originally had is equal to x.
he gave 40% to his wife and 25% to his son, so he gave away 65% of the original amount to his wife and son.

he is left with x - .65*x = .35*x

he spent 60% of the remaining amount on education.

he spent .60 * .35 * x = .21 * x.

he is now left with .35 * x - .21 * x = .14 * x.

that amount is equal to 2700, so you get:

.14 * x = 2700

divide both sides by .14 to get:

x = 2700 / .14 = 19285.71429...

let's see if that works.

he started with 19285.71429...

he gave 40% to his wife and 25% to his son.

he is left with .35 * 19285.71429... which is equal to 6750.

he spent 60% of that on education.

he is left with .4 * 6750 = 2700.

the numbers work out.

he started with 19285.71429... which you can round off to 19285.71 if you need to.